# LLM-based Agents Suffer from Hallucinations: A Survey of Taxonomy, Methods, and Directions

### v1

@@ 3.2. Execution Hallucinations
The execution phase is where the agent translates its deliberated plan $p_{t}$ into a concrete, executable action $a_{t}$. $a_{t}$ typically involves invoking external tools, either as a single operation or through multiple parallel instances (Kim et al., 2024; Singh et al., 2024; Zhu et al., 2025c). Broadly, this execution process in Eq.(2) and Eq.(9) can be further decomposed into two sub-stages (Qu et al., 2025b; Song et al., 2023; Patil et al., 2024; Qin et al., 2024):
@@
    (21) $T_{s}=\text{Select}\left(b_{t},p_{t},\mathcal{T}_{\text{cand}}\right).$
    As there are a large number of tools, $\mathcal{T}_{\text{cand}}$ is typically retrieved from a full tool set $\mathcal{T}$ to narrow the selection scope (Patil et al., 2024; Qin et al., 2024; Gan and Sun, 2025; Qu et al., 2024; Xu et al., 2024a):
    (22) $\mathcal{T}_{\text{cand}}=\text{Retr}\left(b_{t},p_{t},\mathcal{T}\right).$
  • Tool Calling: Once $T_{s}$ is selected, the agent then needs to populate $T_{s}$ with the tool parameters derived from $b_{t}$ and $p_{t}$ to form the final executable action $a_{t}$ (Hsieh et al., 2023; Liu et al., 2024a; Wang et al., 2024c):
    (23) $a_{t}=\mathrm{Call}\left(b_{t},p_{t},T_{s}\right).$
@@ 3.2.1. Tool Documentation Limitation
Execution hallucinations often arise when the agent’s internal representation of tool behaviors diverges from the actual functionality of tools, frequently due to limitation in Tool Documentation (Hsieh et al., 2023). These hallucinations occur because the agent believes it has correctly selected or invoked a tool, even though its decision is based on misleading documentation. Such deficiencies may include redundant information (Yuan et al., 2024b; Qu et al., 2025a), incomplete or inaccurate descriptions (Yuan et al., 2024b; Qu et al., 2025a; Milev et al., 2025; Ni et al., 2025), or a lack of standardization (Yuan et al., 2024b; Shi et al., 2025; Lu et al., 2025; Ni et al., 2025), all of which impair the agent’s ability to properly use tools.
3.2.2. Shallow Pattern Understanding
@@
The agent’s shallow understanding of Tool Patterns may lead to execution hallucinations wherein the agent confidently invokes invalid or outdated tools, mistakenly assuming successful execution. In fact, effective tool usage requires LLM-based agents to develop deep understanding of tool patterns. This requirement primarily stems from two key factors: First, tools often provide multiple functionalities that can be applied across diverse scenarios, rather than being limited to a single task-specific use case (Gao et al., 2024); Second, tools frequently exhibit complex collaborative patterns, such as sequential dependencies and the need for nested calls (Abdelaziz et al., 2024). However, LLM-based agents are typically trained with insufficient exposure to diverse and complex tool-use scenarios (Wang et al., 2024c; Tang et al., 2023a; Zeng et al., 2025; Patil et al., 2024; Qin et al., 2024). This limitation makes agents prone to hallucinating tool invocations that may appear plausible or confident, but in fact violate expected patterns or omit essential details, especially when dealing with complex and novel tasks (Chen et al., 2024k; Cheng and Amiri, 2025; Bai et al., 2024; Han et al., 2025b; He et al., 2025b).
3.2.3. Weak Dynamic Adaptability
@@
LLM-based agents are typically trained on relatively static datasets, with their tool-use knowledge embedded in fixed model parameters (Patil et al., 2024; Qin et al., 2024). However, the execution environment of tools is inherently dynamic and continuously evolving. Tool functionalities may evolve, API interfaces may be modified, and in some cases, tools may be deprecated or replaced (Chen et al., 2024k; Cheng and Amiri, 2025; Bai et al., 2024; Lumer et al., 2025b; Sun et al., 2024; Ding, 2025; Xian et al., 2025). When an agent lacks sufficient adaptability to these Tool Dynamics, its tool-use behavior becomes misaligned with the actual environment, thus leading to execution hallucinations wherein the agent confidently invokes outdated or invalid tools while incorrectly assuming successful execution.
3.2.4. Lack of Solvability Awareness
@@
Tool Solvability refers to whether the current plan $p_{t}$ can be successfully executed under existing conditions. It is primarily related to two factors: (1) the availability of suitable tools in $\mathcal{T}$; (2) the clarity and completeness of $p_{t}$. A lack of solvability awareness in LLM-based agents can also lead to execution hallucinations, where the agent mistakenly assumes that $p_{t}$ is solvable and proceeds with unjustified confidence. For example, if no suitable tools are available, any tool retrieved from $\mathcal{T}$ is likely to be irrelevant or even fabricated and non-executable, resulting in tool selection hallucinations (Zhang et al., 2024b; Xu et al., 2024c). Furthermore, even when a suitable tool exists and is correctly selected, an unclear or incomplete plan $p_{t}$ can prevent the agent from correctly deriving the required parameters for tool invocation (Wang et al., 2025d; Xu et al., 2024c; Wang et al., 2024g; Zhang et al., 2025d; Chen et al., 2025b). Consequently, the agent may omit required parameters or fabricate parameter values, falsely believing that the tool call is valid.
3.3. Perception Hallucinations
@@
Analogous to how humans depend on sensory organs like eyes and ears to acquire information from the external world and convert it into neural signals, LLM-based agents also rely on a perception module to facilitate interaction with the learning environment. The perception module serves as a fundamental interface, extending its perception into a multi-modal space that includes textual, auditory, and visual modalities. As illustrated in Eq.(5) and Eq.(13), the agent receives the external information (e.g., $s_{t+1}$) and transforms this information into internal observations (e.g., $o_{t+1}$) via the perception module. Here we analyze the causes of perception hallucinations.
3.3.1. Environmental Sensor Malfunction
Agents typically rely on environmental sensors to collect external data and convert it into digital signals (Akyildiz et al., 2002). Common sensor types include visual sensors (e.g., cameras), auditory sensors (e.g., microphones), tactile sensors (e.g., pressure-sensitive pads), and inertial measurement units (e.g., accelerometers) (Liu et al., 2025d). When environmental sensor malfunctions occur, such as lens distortion in cameras or signal drift in inertial measurement units, agents may fail to accurately receive raw input information, potentially resulting in perception hallucinations (Trivedi et al., 2025).
@@ 3.3.1. Environmental Sensor Malfunction
3.3.2. Limited Encoding Capability
@@
After receiving $s_{t+1}$, LLM-based agents typically employ a specific encoding module to generate $o_{t+1}$, which serves as a high-level representation of the received signals such as images and sound waves. However, this encoding process encounters two critical limitations: (1) Insufficient Unimodal Representation: Agents struggle to extract the key information of individual modality from $s_{t+1}$. This limitation arises from multiple factors, potentially including insufficient training data quality (Yu et al., 2024a; Liu et al., 2024c), the Transformer architecture’s tendency (Vaswani et al., 2017) to overlook local details (Chen et al., 2024h; Mari and Milani, 2024), and training conflicts between the pre-training and fine-tuning stages (Gekhman et al., 2024; Liu et al., 2025f). (2) Weak Cross-modal Collaboration: Agents lack an effective mechanism to integrate semantic associations across different modalities, resulting in the incomplete representation of $s_{t+1}$. When faced with a complex learning environment, LLM-based agents typically perform the procedures of modality alignment and joint encoding to integrate different multi-modal information (Ghazanfari et al., 2024; Wang et al., 2024a; Zhao et al., 2025b). However, during this process, LLM-based agents may fail to align modalities due to insufficient cross-modal annotations and imbalanced modality data (Sun et al., 2023; Chen et al., 2023b; Lai et al., 2024a). Furthermore, in the generation stage, agents may be dominated by the language prior, gradually neglecting non-textual information (Huang et al., 2024a; Wang et al., 2024h).
3.4. Memorization Hallucinations
The memory module is a core component of LLM-based agents, tasked with storing and managing information to facilitate subsequent decision-making (Xiong et al., 2025a; Lu et al., 2023; Packer et al., 2023; Huang et al., 2023). It primarily owns two operations: Memory Retrieval: Extracting and integrating relevant information from stored memory to support the current decision process; Memory Update: revising and removing existing memory based on newly acquired information or feedback to ensure its accuracy and timeliness.
@@ 3.4. Memorization Hallucinations
Definition of Memorization Hallucinations Memorization hallucinations refer to the phenomenon in which an agent implicitly assumes its memory to be accurate and reliable, without validating the correctness of the stored content. Consequently, when the agent incorporates its memory module into decision-making, it may rely on outdated, fabricated, or conflated memories, resulting in flawed reasoning or inappropriate actions. Based on the above two operations, we categorize memorization hallucinations into two types: Memory Retrieval Hallucinations, where irrelevant or non-existent information are retrieved; and Memory Update Hallucinations, involving incorrect modifications and deletions of memory content.
@@ 3.4.1. Flawed Memory Initialization
The Initial Memory $m_{0}$ serves as the foundation of memorization in LLM-based agents, and its quality directly influences the reliability of subsequent memory processes. Biased or incomplete content in $m_{0}$ can introduce initialization deficiencies, leading to memorization hallucinations before deployment. In particular, certain biases, particularly those related to gender and nationality, are inherently linked to hallucination issues (Huang et al., 2025d). Moreover, when critical information is missing (e.g., the time or location of a historical event) in $m_{0}$, agents may rely on inherent biases or assumptions to fill in the gaps, resulting in inaccurate or fabricated responses in many real-world scenarios (Li et al., 2024b; Singh et al., 2025; Singhal et al., 2025; Kasai et al., 2023).
3.4.2. Sub-optimal Retrieval Mechanism
@@
First, poor Ranking Strategies may lead the agent to retrieve memory content that only superficially ”appears similar” but lacks true relevance (Tu et al., 2025; Agrawal et al., 2024; Barnett et al., 2024). Additionally, deficiencies in the memory Indexing Structure (He et al., 2024; Barnett et al., 2024; Modarressi et al., 2023), such as inappropriate index granularity or delayed index updates, can further exacerbate this problem, potentially resulting in information loss and the retrieval of outdated memory. Finally, insufficient understanding of Query Semantics is also an important problem. If the agent misinterprets the query’s intent or overlooks specific task requirements (Peng et al., 2023; Zhang and Zhang, 2025; Chen et al., 2024a; Yoon et al., 2025b), it may fail to capture contextual cues necessary for accurate retrieval.
3.4.3. Imperfect Priority Assignment
@@
This issue mainly manifests in two aspects: Forgetting Priority and Merging Priority. For memory forgetting, poorly assigned priorities can result in the elimination of important information or the retention of irrelevant content, ultimately compromising the accuracy of subsequent decisions (Zhong et al., 2024; Packer et al., 2023; Mei et al., 2025). Moreover, when the agent merges multiple memory fragments, failure to correctly assess their priorities (Jin et al., 2024a; Ding et al., 2024) may result in the merged memory containing inherent conflicts (Wang et al., 2025i; Wu et al., 2024b; Xie et al., 2023). Some of these conflicts are implicit (Xu et al., 2024b): Although memory fragments may appear similar, they differ significantly in key semantic details, thereby increasing the difficulty of priority determination.
3.4.4. Inaccurate Information Writing
@@
When performing long-term tasks and engaging in multi-turn interactions, agents must summarize, structure, and store relevant historical information to the memory module. However, this process is susceptible to Information Compression issues, where the generated summaries may be overly general, omit crucial details (Xu et al., 2025c; Maharana et al., 2024; Qin et al., 2025; Wang et al., 2025b), or introduce distortions due to imperfect abstraction (Guo et al., 2025a; Ye, 2025; Shinn et al., 2023; Xiong et al., 2025a). Additionally, non-standardized Memory Formats with disorganized structures can hinder writing efficiency (Xiong et al., 2025a; Wang and Chen, 2025). Memory Capacity Constraints exacerbate these challenges, because limited storage may necessitate the selective retention of information (Packer et al., 2023; Zhang et al., 2024a), increasing the risk that salient but less frequently accessed information is discarded.
3.5. Communication Hallucinations
Compared with the independent operations of a single agent, LLM-based MAS emphasizes collaboration and coordination to harness collective intelligence for solving more complex and demanding tasks (Ha and Tang, 2022; Li et al., 2023a; Tran et al., 2025). In such systems, effective inter-agent communication serves as an essential requirement (Yan et al., 2025; Li et al., 2024c), facilitating the exchange of ideas and the coordination of plans among agents (Piatti et al., 2024; Wang et al., 2025c; de Cerqueira et al., 2024).
@@ 3.5. Communication Hallucinations
Communication Hallucinations in LLM-based Agents Communication hallucinations refer to the phenomenon in which LLM-based agents appear to engage in meaningful inter-agent communication, yet the information exchanged is inaccurate, misleading, or fabricated, thereby undermining collaboration. Such hallucinations typically arise from Erroneous Message Propagation, where incorrect information is shared as if reliable; Uncoordinated Communication Protocols, where inconsistent formats or timing lead to misunderstandings; and Ineffective Network Updates, where outdated or poorly synchronized connections distort information flow and hinder task execution.
@@ 3.5.1. Erroneous Message Propagation
In LLM-based MAS, agents fundamentally rely on LLMs (Xi et al., 2025; Cheng et al., 2024b; Liu et al., 2025d) to generate messages for exchanging information with other agents. However, since LLMs are prone to the well-known Factuality and Faithfulness Hallucinations (Zhang et al., 2023b; Huang et al., 2025d), some agents may produce messages containing inaccurate facts, misinterpretations of shared knowledge, or misleading inferences (Yoffe et al., 2024; Cemri et al., 2025; Zhang et al., 2025g), thereby giving rise to communication hallucinations. Beyond this, Content Redundancy is also an important cause, where agents generate unnecessary or repetitive content that obscures critical signals, increases cognitive load, and sometimes leads to redundant task execution steps that manifest as logical errors (Zhang et al., 2025h; Wang et al., 2025f; Guo et al., 2024b; Cemri et al., 2025). Information Asymmetry can further exacerbate this issue. Because agents own different roles and positions within MAS, the information accessible to them varies, and such asymmetric settings may yield vague or incomplete instructions that hinder task comprehension (Liu et al., 2024b) and amplify the risk of biased decisions (Wang et al., 2025a; Zhang et al., 2025e).
3.5.2. Uncoordinated Communication Protocols
@@
Communication protocols govern how agents exchange messages, directly determining the efficiency, reliability, and coordination of their interactions (Ehtesham et al., 2025; Yan et al., 2025; Yang et al., 2025c). Without a unified and effective protocol, agents may ”talk past each other”, leading to communication hallucinations. First, LLM-based MAS usually follows a manner of Asynchronous Scheduling (Hong et al., 2024), so when receiving and processing instructions, agents may encounter issues of information loss and information overload. Such temporal discrepancies can result in information errors, thereby increasing the risk of hallucinatory outputs (Gim et al., 2024; Gonzalez-Pumariega et al., 2025). Second, communication protocols define the Message formats. Current LLM-based agents predominantly rely on the format of natural language (Bansal et al., 2024) which often introduces instruction ambiguity. Adopting structured formats (e.g., JSON) can improve clarity and rigor of expression, which mitigates the risk of miscommunication (Yang et al., 2024; Wang et al., 2025e). Finally, LLM-based MAS demands a robust Fault-tolerant Design, incorporating confirmation conditions and synchronization constraints to avoid erroneous decisions caused by message loss or delays in dynamic or noisy environments (Liu et al., 2024d; Kwon and Zhang, 2025).
3.5.3. Ineffective Network Updates
@@
Network topology defines how agents are interconnected, determining who communicates with whom and how frequently (Yan et al., 2025; Park et al., 2023; Chen et al., 2024f). As discussed in Section 2.3, the network topology is continuously evolving, and network updates reshape the propagation paths of messages within MAS. When network updates are ineffective, they can induce communication hallucinations due to inconsistent or outdated inter-agent connections (Cemri et al., 2025; Zhang et al., 2025g; Zhuge et al., 2024). Although recently proposed strategies improve the flexibility and responsiveness of MAS, they often suffer from delayed updates and poor coordination (Park et al., 2023; Song et al., 2025; Aratchige and Ilmini, 2025). If the updated network fails to accurately reflect agents’ current relevance or expertise, messages may be routed to inappropriate recipients, leading to misunderstandings or redundant reasoning in LLM-based MAS.
4. Systematic Methodology

### v2

@@ III-B Execution Hallucinations
The execution phase is where the agent translates its deliberated plan $p_{t}$ into a concrete, executable action $a_{t}$. $a_{t}$ typically involves invoking external tools, either as a single operation or through multiple parallel instances [zhu2025divide]. Broadly, this execution process in Eq.(2) can be further decomposed into two sub-stages [patil2024gorilla]:
@@
    $T_{s}=\text{Select}\left(b_{t},p_{t},\mathcal{T}_{\text{cand}}\right).$ (12)
    As there are a large number of tools, $\mathcal{T}_{\text{cand}}$ is typically retrieved from a full tool set $\mathcal{T}$ to narrow the selection scope [qin2023toolllm]:
    $\mathcal{T}_{\text{cand}}=\text{Retr}\left(b_{t},p_{t},\mathcal{T}\right).$ (13)
  • Tool Calling: Once $T_{s}$ is selected, the agent then needs to populate $T_{s}$ with the tool parameters derived from $b_{t}$ and $p_{t}$ to form the final executable action $a_{t}$:
    $a_{t}=\mathrm{Call}\left(b_{t},p_{t},T_{s}\right).$ (14)
@@
1) Tool Documentation Limitation. Execution hallucinations often arise when the agent’s internal representation of tool behaviors diverges from the actual functionality of tools, frequently due to limitation in Tool Documentation [r32]. These hallucinations occur because the agent believes it has correctly selected or invoked a tool, even though its decision is based on misleading documentation. Such deficiencies may include redundant information [r12, qu2025from], incomplete or inaccurate descriptions [r13, r16], or a lack of standardization [r14, r15], all of which impair the agent’s ability to properly use tools.
@@
2) Shallow Pattern Understanding. The agent’s shallow understanding of Tool Patterns may lead to execution hallucinations wherein the agent confidently invokes invalid or outdated tools, mistakenly assuming successful execution. In fact, effective tool usage requires LLM-based agents to develop deep understanding of tool patterns. This requirement primarily stems from two key factors: First, tools often provide multiple functionalities that can be applied across diverse scenarios, rather than being limited to a single task-specific use case [r33]; Second, tools frequently exhibit complex collaborative patterns, such as sequential dependencies and the need for nested calls [r311]. However, LLM-based agents are typically trained with insufficient exposure to diverse and complex tool-use scenarios [r34, r35, r36]. This limitation makes agents prone to hallucinating tool invocations that may appear plausible or confident, but in fact violate expected patterns or omit essential details, especially when dealing with complex and novel tasks [r22].
@@
3) Weak Dynamic Adaptability. LLM-based agents are typically trained on relatively static datasets, with their tool-use knowledge embedded in fixed model parameters [r35, r36]. However, the execution environment of tools is inherently dynamic and continuously evolving. Tool functionalities may evolve, API interfaces may be modified, and in some cases, tools may be deprecated or replaced [r22, lumer2025scalemcp, xian2025measuring]. When an agent lacks sufficient adaptability to these Tool Dynamics, its tool-use behavior becomes misaligned with the actual environment, thus leading to execution hallucinations wherein the agent confidently invokes outdated or invalid tools while incorrectly assuming successful execution.
@@
4) Lack of Solvability Awareness. Tool Solvability refers to whether the current plan $p_{t}$ can be successfully executed under existing conditions. It is primarily related to two factors: a) the availability of suitable tools in $\mathcal{T}$; b) the clarity and completeness of $p_{t}$. A lack of solvability awareness in LLM-based agents can also lead to execution hallucinations, where the agent mistakenly assumes that $p_{t}$ is solvable and proceeds with unjustified confidence. For example, if no suitable tools are available, any tool retrieved from $\mathcal{T}$ is likely to be irrelevant or even fabricated and non-executable, resulting in tool selection hallucinations [r54]. Furthermore, even when a suitable tool exists and is correctly selected, an unclear or incomplete plan $p_{t}$ can prevent the agent from correctly deriving the required parameters for tool invocation [wang2024learning]. Consequently, the agent may omit required parameters or fabricate parameter values, falsely believing that the tool call is valid.
III-C Perception Hallucinations
@@
Analogous to how humans depend on sensory organs like eyes and ears to acquire information from the external world and convert it into neural signals, LLM-based agents also rely on a perception module to facilitate interaction with the learning environment. The perception module serves as a fundamental interface, extending its perception into a multi-modal space that includes textual, auditory, and visual modalities. As illustrated in Eq.(5), the agent receives the external information (e.g., $s_{t+1}$) and transforms this information into internal observations (e.g., $o_{t+1}$) via the perception module. Here we analyze the causes of perception hallucinations.
@@
1) Environmental Sensor Malfunction. Agents typically rely on environmental sensors to collect external data and convert it into digital signals [trivedi2025intelligent]. Common sensor types include visual sensors (e.g., cameras), auditory sensors (e.g., microphones), tactile sensors (e.g., pressure-sensitive pads), and inertial measurement units (e.g., accelerometers) [liu2025advances]. When environmental sensor malfunctions occur, such as lens distortion in cameras or signal drift in inertial measurement units, agents may fail to accurately receive raw input information, potentially resulting in perception hallucinations.
@@
2) Limited Encoding Capability. After receiving $s_{t+1}$, LLM-based agents typically employ a specific encoding module to generate $o_{t+1}$, which serves as a high-level representation of the received signals such as images and sound waves. However, this encoding process encounters two critical limitations: a) Insufficient Unimodal Representation: Agents struggle to extract the key information of individual modality from $s_{t+1}$. This limitation arises from multiple factors, potentially including insufficient training data quality [yu2024hallucidoctor], the Transformer architecture’s tendency [vaswani2017attention] to overlook local details [chen2024florence], and training conflicts between the pre-training and fine-tuning stages [liu2025visual]. b) Weak Cross-modal Collaboration: Agents lack an effective mechanism to integrate semantic associations across different modalities, resulting in the incomplete representation of $s_{t+1}$. When faced with a complex learning environment, LLM-based agents typically perform the procedures of modality alignment and joint encoding to integrate different multi-modal information [ghazanfari2024emma]. However, during this process, LLM-based agents may fail to align modalities due to insufficient cross-modal annotations and imbalanced modality data [lai2024lisa]. Furthermore, in the generation stage, agents may be dominated by the language prior, gradually neglecting non-textual information [wang2024vigc].
@@
Definition of Memorization Hallucinations Memorization hallucinations refer to the phenomenon in which an agent implicitly assumes its memory to be accurate and reliable, without validating the correctness of the stored content. Consequently, when the agent incorporates its memory module into decision-making, it may rely on outdated, fabricated, or conflated memories, resulting in flawed reasoning or inappropriate actions. Based on the above two operations, we categorize memorization hallucinations into two types: Memory Retrieval Hallucinations (MRHs), where irrelevant or non-existent information are retrieved; and Memory Update Hallucinations (MUHs), involving incorrect modifications and deletions of memory content.
III-D Memorization Hallucinations
The memory module is a core component of LLM-based agents, tasked with storing and managing information to facilitate subsequent decision-making [xiong2025memorymanagementimpactsllm, packer2023memgpt, huang2023memory]. It primarily owns two operations: Memory Retrieval: Extracting and integrating relevant information from stored memory to support the current decision process; Memory Update: revising and removing existing memory based on newly acquired information or feedback to ensure its accuracy and timeliness.
@@ III-D Memorization Hallucinations
1) Flawed Memory Initialization. The Initial Memory $m_{0}$ serves as the foundation of memorization in LLM-based agents, and its quality directly influences the reliability of subsequent memory processes. Biased or incomplete content in $m_{0}$ can introduce initialization deficiencies, leading to memorization hallucinations before deployment. In particular, certain biases, particularly those related to gender and nationality, are inherently linked to hallucination issues [huang2025survey]. Moreover, when critical information is missing (e.g., the time or location of a historical event) in $m_{0}$, agents may rely on inherent biases or assumptions to fill in the gaps, resulting in inaccurate or fabricated responses in many real-world scenarios [li2024banishing, singh2025agentic, singhal2025toward].
@@
2) Sub-optimal Retrieval Mechanism. First, poor Ranking Strategies may lead the agent to retrieve memory content that only superficially ”appears similar” but lacks true relevance [tu2025rbft, Agrawal2024MindfulRAGAS]. Additionally, deficiencies in the memory Indexing Structure [he2024g, Modarressi2023RETLLMTA], such as inappropriate index granularity or delayed index updates, can further exacerbate this problem, potentially resulting in information loss and the retrieval of outdated memory. Finally, insufficient understanding of Query Semantics is also an important problem. If the agent misinterprets the query’s intent or overlooks specific task requirements [zhang2025hallucination, yoon2025ask], it may fail to capture contextual cues necessary for accurate retrieval.
@@
3) Imperfect Priority Assignment. This issue mainly manifests in two aspects: Forgetting Priority and Merging Priority. For memory forgetting, poorly assigned priorities can result in the elimination of important information or the retention of irrelevant content, ultimately compromising the accuracy of subsequent decisions [packer2023memgpt, mei2025survey]. Moreover, when the agent merges multiple memory fragments, failure to correctly assess their priorities [jin2024tug, ding2024retrieve] may result in the merged memory containing inherent conflicts [wang2025accommodate, wu2024faithful]. Some of these conflicts are implicit [xu2024knowledge]: Although memory fragments may appear similar, they differ significantly in key semantic details, thereby increasing the difficulty of priority determination.
@@
4) Inaccurate Information Writing. When performing long-term tasks and engaging in multi-turn interactions, agents must summarize, structure, and store relevant historical information to the memory module. However, this process is susceptible to Information Compression issues, where the generated summaries may be overly general, omit crucial details [xu2025mem, qin2025towards, wang2025recursively], or introduce distortions due to imperfect abstraction [Guo2025RepoAuditAA, ye2025task, xiong2025memorymanagementimpactsllm]. Additionally, non-standardized Memory Formats with disorganized structures can hinder writing efficiency [xiong2025memorymanagementimpactsllm, wang2025mirix]. Memory Capacity Constraints exacerbate these challenges, because limited storage may necessitate the selective retention of information [packer2023memgpt, zhang2024survey], increasing the risk that salient but less frequently accessed information is discarded.
III-E Communication Hallucinations
@@
Definition of Communication Hallucinations Communication hallucinations refer to the phenomenon in which LLM-based agents appear to engage in meaningful inter-agent communication, yet the information exchanged is inaccurate, misleading, or fabricated, thereby undermining collaboration. Such hallucinations typically arise from erroneous message propagation, where incorrect information is shared as if reliable; uncoordinated communication protocols, where inconsistent formats or timing lead to misunderstandings; and ineffective network updates, where outdated or poorly synchronized connections distort information flow and hinder task execution.
Compared with the independent operations of a single agent, LLM-based MAS emphasizes collaboration and coordination to harness collective intelligence for solving more complex and demanding tasks [tran2025multi]. In such systems, effective inter-agent communication serves as an essential requirement [yan2025beyond], facilitating the exchange of ideas and the coordination of plans among agents [piatti2024cooperate].
@@
1) Erroneous Message Propagation. In LLM-based MAS, agents fundamentally rely on LLMs [xi2025rise, cheng2024exploring, liu2025advances] to generate messages for exchanging information with other agents. However, since LLMs are prone to the well-known Factuality and Faithfulness Hallucinations [huang2025survey], some agents may produce messages containing inaccurate facts, misinterpretations of shared knowledge, or misleading inferences [yoffe2024debunc, cemri2025multi, zhang2025agent], thereby giving rise to communication hallucinations. Beyond this, Content Redundancy is also an important cause, where agents generate unnecessary or repetitive content that obscures critical signals, increases cognitive load, and sometimes leads to redundant task execution steps that manifest as logical errors [zhang2025cut, wang2025agentdropout, cemri2025multi]. Information Asymmetry can further exacerbate this issue. Because agents own different roles and positions within MAS, the information accessible to them varies, and such asymmetric settings may yield vague or incomplete instructions that hinder task comprehension [liu2024autonomous] and amplify the risk of biased decisions [wang2025learning].
@@
2) Uncoordinated Communication Protocols. Communication protocols govern how agents exchange messages, directly determining the efficiency, reliability, and coordination of their interactions [ehtesham2025survey, yan2025beyond, yang2025survey]. Without a unified and effective protocol, agents may ”talk past each other”, leading to communication hallucinations. First, LLM-based MAS usually follows a manner of Asynchronous Scheduling [hong2024metagpt], so when receiving and processing instructions, agents may encounter issues of information loss and information overload. Such temporal discrepancies can result in information errors, thereby increasing the risk of hallucinatory outputs [gim2024asynchronous, gonzalez-pumariega2025robotouille]. Second, communication protocols define the Message formats. Current LLM-based agents predominantly rely on the format of natural language [bansal2024challenges] which often introduces instruction ambiguity. Adopting structured formats (e.g., JSON) can improve clarity and rigor of expression, which mitigates the risk of miscommunication [yang2024multi, wang2025talk]. Finally, LLM-based MAS demands a robust Fault-tolerant Design, incorporating confirmation conditions and synchronization constraints to avoid erroneous decisions caused by message loss or delays in dynamic or noisy environments [liu2024a, kwon2025cp].
@@
3) Ineffective Network Updates. Network topology defines how agents are interconnected, determining who communicates with whom and how frequently [yan2025beyond, park2023generative, chen2024agentverse]. As discussed in Appendix A, the network topology is continuously evolving, and network updates reshape the propagation paths of messages within MAS. When network updates are ineffective, they can induce communication hallucinations due to inconsistent or outdated inter-agent connections [cemri2025multi, zhang2025agent, zhuge2024gptswarm]. Although recently proposed strategies improve the flexibility and responsiveness of MAS, they often suffer from delayed updates and poor coordination [park2023generative, song2025code, aratchige2025llms]. If the updated network fails to accurately reflect agents’ current relevance or expertise, messages may be routed to inappropriate recipients, leading to misunderstandings or redundant reasoning in LLM-based MAS.
IV Systematic Methodology